38-) Assume an international regime of fixed exchange rates. When countries do not want to engage in direct intervention, countries with a BOP ________ should consider ________ their currency while countries with a BOP ________ should consider ________ their currency.

Select one:

a. surplus, revaluing; deficit, devaluing

b. surplus, devaluing; deficit, revaluing

c. deficit, devaluing; surplus, devaluing

d. deficit, revaluing; surplus, revaluing

39-)

An MNE has a contract for a relatively predictable long-term inflow of Japanese yen that the firm chooses to hedge by seeking out potential suppliers in Japan. This hedging strategy is referred to as:

Select one:

a. a natural hedge.

b. matching.

c. currency-switching.

d. diversification.

40-)

Assume the current U.S. dollar-yen spot rate is 90 ¥/$. Further, the current nominal 180-day rate of return in Japan is 1% and 2% in the United States. What is the approximate forward exchange rate for 180 days?

Select one:

a. ¥89.55/$

b. ¥90.89/$

c. ¥90.45/$

d. ¥89.12/$
Answer #1

38 Option a.. surplus, revaluing; deficit, devaluing

39 Option a, a. a natural hedge.

40 a .   ¥89.55/$

as per PPP theory, spot rate / future rate = spot currency / forward

[(1+ 2%) / (1+1%)]180/360 = 90 / X

X = 90 * [(1+0.01)/ (1+0.02) ]]180/360

X = 90 * [ 1.01/1.02 ] 0.5

X = 90 * 0.9950

= 89.55